# Machine learning of the Ising model on a spherical Fibonacci lattice

### v1

@@ Appendix A Methodology of phase classification: Details of GCN
$\bm{H}=\operatorname{ReLu}(\bm{L}^{rm}\bm{X}\bm{W}_{h}+\bm{b}_{h})$ (3)
Here, ReLu serves as the activation function, $\bm{W}_{h}\in\mathbb{R}^{1\times 1}$ is the weight, and $\bm{b}_{h}\in\mathbb{R}^{N\times 1}$ is the bias. Finally, we employ a fully connected layer along with the softmax function to aggregate the hidden layer, resulting in an output $\bm{H}\in\mathbb{R}^{N\times 1}$ that generates classification confidence for the ordered and disordered phases, denoted as $p_{\text{o}}$ and $p_{\text{d}}$. The temperature at which $p_{\text{o}}=p_{\text{d}}$ defines the phase transition temperature $T_{c}$.

### v2

@@ Appendix C Methodology of phase classification: Details of GCN
$\bm{H}=\operatorname{ReLu}\left(\bm{L}^{rm}\bm{X}\bm{W}_{h}+\bm{b}_{h}\right)$ (7)
Here, $\operatorname{ReLu}$ serves as the activation function, $\bm{W}_{h}\in\mathbb{R}^{1\times 1}$ is the weight, and $\bm{b}_{h}\in\mathbb{R}^{N\times 1}$ is the bias. Finally, we employ a fully connected layer along with the softmax function to aggregate the hidden layer, resulting in an output $\bm{H}\in\mathbb{R}^{N\times 1}$ that generates classification confidence for the ordered and disordered phases, denoted as $p_{\text{o}}$ and $p_{\text{d}}$. The temperature at which $p_{\text{o}}=p_{\text{d}}$ defines the phase transition temperature $T_{c}$.